OpenStudy (anonymous):

simplify to a single fraction: ((x/x-1)+1)/((x+2)/x)

OpenStudy (anonymous):

\[\left(\frac{x}{x+1}+1 \right)\div \frac{(x+2)}{x}\] this?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

(2x + 1)/(x+1) times x / (x + 2)

OpenStudy (anonymous):

x(2x+1) / (x+1)(x+2)

OpenStudy (anonymous):

wait.... am i multiplying (2x + 1)/(x+1) times x / (x + 2) to the original fraction??

OpenStudy (anonymous):

yes (2x + 1)/(x + 1) is simplification of (1/(x+1) + 1) and since you're dividing by x / x + 2 you multiply by x + 2 / 2

OpenStudy (anonymous):

do you get it?

OpenStudy (anonymous):

soo it would be ((x/x-1)+1)/((x+2)/x) * (2x + 1)/(x+1) times x / (x + 2) ??

OpenStudy (anonymous):

no where you have the * sign should be \[= equal\] to: The expressions are equivalent but in simplified form

OpenStudy (anonymous):

((x/x-1)+1)/((x+2)/x) = (2x + 1)/(x+1) times x / (x + 2)

OpenStudy (anonymous):

[x(2x+1)]/[(x+1)(x+2)] is your answer

OpenStudy (anonymous):

ok i see... and then i would cancel out the x+2 and x correct?

OpenStudy (anonymous):

No, you can only cancel out when you have same terms, x + 2 and x are different terms so you can't cancel them out. That answer will be as simplified as you can get the expression

OpenStudy (anonymous):

so then what would i cancel out in order to simplify it?

OpenStudy (anonymous):

Well in this case you can't cancel anything out it is already simplified tot he lowest form: [x(2x+1)]/[(x+1)(x+2)]

OpenStudy (anonymous):

oh i see.. and what did u do in order to get that? sorry if im asking so many questions.... i just wanna understand it.

OpenStudy (anonymous):

Okay that's fine when you have an algebraic expression like that it is imperative to simplify both before you perform any arithmetic operation..so this was your question ((x/x-1)+1)/((x+2)/x) and we have to simplify this to a single fraction.

OpenStudy (anonymous):

First of all we take : [(x/(x-1) + 1] = x/(x-1) + (x - 1)/(x - 1) = 2x - 1 / x -1..right?

OpenStudy (anonymous):

right...

OpenStudy (anonymous):

So know we have [(2x-1)/(x-1)] / [(x+2)/x] [(2x-1)/(x-1)] x [x/(x+2)]

OpenStudy (anonymous):

Answer = [x(2x-1)]/[(x-1)(x+2)]

OpenStudy (anonymous):

OH! i get it now. thank you!!

OpenStudy (anonymous):

Great! You're welcome!